OpenStudy (anonymous):

5x+9y=8 -8x+y=18 please do not answer unless you are for sure if i fail this class i dont get to graduate college, to me my degree means everything please someone help me who cares

OpenStudy (anonymous):

(-2,2)

OpenStudy (anonymous):

promise :)

hero (hero):

There's no way you're computing those values that fast.... What program are you using?

OpenStudy (anonymous):

ty you are correct omg i got 5 more questions dont go away please i need to pass im so nervous im crying

OpenStudy (anonymous):

my head....its simple?....

hero (hero):

FYI, we didn't get the first one wrong either...

OpenStudy (turingtest):

i agree

hero (hero):

I think you posted something wrong. Two different people can't compute the same wrong answer, sorry.

hero (hero):

Especially when it results in a fraction.

OpenStudy (anonymous):

4x+5y=2 -4x+y=34

hero (hero):

Let's see how fast QED posts this one

OpenStudy (anonymous):

i was answering another dude sorry to disappoint you hero :P

OpenStudy (anonymous):

QED is smart dont hate apprectiate ignore them just help me lol

OpenStudy (turingtest):

6y=36 y=6 x=7

OpenStudy (anonymous):

turing's answer is right

hero (hero):

yeagerbarbie, don't insult the rest of us because you're stressed

OpenStudy (anonymous):

kidding

OpenStudy (anonymous):

(6,7) is incorrect

OpenStudy (turingtest):

sorry x=-7

OpenStudy (anonymous):

x=-7

OpenStudy (anonymous):

yup

OpenStudy (turingtest):

im using a laptop

OpenStudy (anonymous):

(-7,6)

OpenStudy (anonymous):

(6,-7) is incorrect ok QED ill try your way, you seem to understand this, im going to move you in my home until im done lol

OpenStudy (turingtest):

no (-7,6) is right

hero (hero):

yeager, I haven't gotten any of your questions wrong either. I'm 1-0 so far.

OpenStudy (anonymous):

yay thank you again ok guys 4 more to go, ummm QED only please that person knows my math thank you so very much btw!!!

OpenStudy (anonymous):

hero sorry didnt man to offend you im in college classes and can not afford to fail i do again apologize

hero (hero):

He's just really fast. I answered your first question correctly. QED and I got the exact same answer. I haven't answered any of your other ones yet.

OpenStudy (anonymous):

next question 5x+2y=-13 7x-3y=19

OpenStudy (turingtest):

yes we can all do these here and we just wanted to beat qed, who accepted my wrong answer as correct btw

OpenStudy (anonymous):

sprry again Hero i di not mean to offend anyone

OpenStudy (anonymous):

fast bcuz im a math major woot woot :) go McGill !

OpenStudy (anonymous):

ok QED whats it going to take to move you closer to teach me mr math wizard...I got final exam in 2 weeks form sunday, i go to online so mine are on sundays...

hero (hero):

(-1,-4)

hero (hero):

Try that

OpenStudy (anonymous):

sorry hero that answer is incorrect i have 2 more trys here :(

OpenStudy (turingtest):

we must check our answers, these seem important to her

hero (hero):

yeager, maybe if you didn't panic so much, I wouldn't either

OpenStudy (anonymous):

(-1/29, -186/29)

OpenStudy (anonymous):

Yeagear, are you sure you put the parentheses around your answer?

OpenStudy (anonymous):

my answer is right yeager

OpenStudy (turingtest):

yes he is right

hero (hero):

I subtracted 14 from both sides instead of adding 14. My bad. I fail under pressure. I would have been more careful, but I was trying to be superman like Outkast

OpenStudy (anonymous):

You ARE superman, my Hero.